# IncomeSCM: From tabular data set to time-series simulator and causal estimation benchmark

### v1

@@ 5 Experiments
We see a generally good match in the base statistics of variables, with the biggest discrepancy for income>50k. This is due to renormalizing to achieve an average income of \$70 000 before applying noise and correction for the effects of studies and workclass = Without-pay. The discriminative performance for the fitted samplers varies greatly between variables, some (such as marital-status being more predictable from their parents in the posited causal graph than others (such as capital-net). This is expected since the original data set was not constructed with predicting all variables in mind, nor are all available variables used for prediction, only the causal parents. However, it does mean that the values of variables with low discriminative performance are largely determined by exogenous noise.
Table 2: Results estimating CATE and ATE using the full IncomeSCM-1.0.CATE data set. $R^{2}$ is the coefficient of determination, AE is the absolute error. AUC/$R^{2}$ CV are the 5-fold cross-validation AUC of the propensity estimate, and $R^{2}$ of the observed outcome, for propensity and regression estimators, respectively. For $R^{2}$ CATE and AE ATE, the confidence intervals ($\alpha=0.05$) are computed by 1000 bootstrap iterations over test samples. An unbiased estimate of the ATE is \$40919.
@@
Adjustment set: All covariates
IPW (LR) -0.16 (-0.16, -0.15) 36855 (36081, 37676) 0.94
IPW (RF) -0.75 (-0.77, -0.74) 81120 (80348, 81905) 0.94
IPW-W (LR) -0.00 (-0.00, -0.00) 3992 (3161, 4841) 0.94
IPW-W (RF) -0.01 (-0.01, -0.01) 10511 (9683, 11317) 0.94
S-learner (Ridge) -0.04 (-0.05, -0.04) 19140 (18382, 19936) 0.53
S-learner (XGB) 0.22 (0.22, 0.23) 747 (97, 1487) 0.60
S-learner (RF) 0.19 (0.18, 0.19) 3235 (2464, 4001) 0.61
T-learner (Ridge) 0.23 (0.22, 0.24) 2029 (1288, 2770) 0.57
T-learner (XGB) 0.23 (0.23, 0.24) 2599 (1865, 3307) 0.60
T-learner (RF) 0.26 (0.25, 0.27) 3646 (2945, 4334) 0.61
Adjustment set: Direct causes of treatment only
IPW (LR) -0.00 (-0.00, -0.00) 2338 (1564, 3158) 0.91
IPW (RF) -0.64 (-0.66, -0.63) 74985 (74213, 75771) 0.93
IPW-W (LR) -0.00 (-0.00, -0.00) 3045 (2196, 3876) 0.91
IPW-W (RF) -0.01 (-0.01, -0.01) 9360 (8532, 10165) 0.93
S-learner (Ridge) -0.04 (-0.05, -0.04) 19425 (18667, 20221) 0.52
S-learner (XGB) 0.22 (0.21, 0.22) 2513 (1774, 3206) 0.57
S-learner (RF) 0.15 (0.14, 0.15) 3399 (2616, 4186) 0.56
T-learner (Ridge) 0.21 (0.20, 0.22) 6659 (5902, 7413) 0.56
T-learner (XGB) 0.15 (0.14, 0.15) 3045 (2246, 3837) 0.56
T-learner (RF) 0.23 (0.22, 0.24) 2678 (1967, 3417) 0.57
@@
Regression adjustment estimators directly model the outcome $Y$ as a function of the intervention $A$ and adjustment variables $Z$ which guarantee identifiability. Among these, we include examples of T-learners, which fit one model per potential outcome (intervention), and S-learners, which treat the intervention as an input to a single model, mixed in with adjustment variables (Künzel et al., 2019). For both so-called “meta learners”, we compare multiple base estimators in ridge regression (Ridge), random forests (RF), and XGBoost (XGB). Inverse propensity-weighting estimators model the intervention variable instead, as a function of the adjustment set, and re-weigh outcomes to estimate treatment effects. We use two classical estimators, the Horvitz-Thompson (IPW) and Hayek or “Weighted” (IPW-W) estimators of the average treatment effect (ATE). These return only the ATE, not the CATE. For Tasks 1–2, all estimators return estimates of the CATE conditioned on the given set of variables and adjusting for that same set of variables. In the case of regression estimators returning prediction models $\hat{\mu}_{a}(z)\approx\mathbb{E}[Y(a)\mid Z=z]$ of the potential outcome for each intervention, $a\in\{0,1\}$ and a conditioning set $Z$, $\hat{\mbox{CATE}}(z)=\hat{\mu}_{1}(z)-\hat{\mu}_{0}(z).$ The ATE is estimated as an average of $\hat{\mbox{CATE}}(z)$ across the evaluation set. For IPW estimators, a propensity model $\hat{e}(z)\approx p(T=1\mid Z=z)$ is used to compute an estimate of each expected potential outcome $\hat{\mu}_{a}\approx\mathbb{E}[Y(a)]$ which is used in place of $\mu_{a}(z)$ in $\hat{\mbox{CATE}}(z)$. For Task 3, the CATE is estimated as in Task 2 but stratified and aggregated in 10-year bins of hours-per-week. Ground truth for each task is constructed analogously but with sampled potential outcomes in place of $\hat{\mu}$.
Model selection.
Hyperparameters for all estimators were selected based on 5-fold cross-validation of 20 uniformly sampled hyperparameter settings. The estimators were then refitted to the entirety of the observational training data. Ranges for hyperparameters are specified in Appendix Table 6. Producing all results, including fitting the simulator, sampling, estimating and evaluating causal effects, took less than 1 hour on a 2023 M2 MacBook Pro.
@@ Results.
We present the results for causal effect estimation Tasks 1–2 in Table 2 and for Task 3 in Table 3 and Figure 2. IPW methods and S-learner (Ridge) are excluded from the latter two since these methods do not yield heterogenous effect estimates, only ATE. All tasks are challenging to solve: the best methods achieve CATE $R^{2}$ of only 0.26, 0.23 and 0.59 on Tasks 1, 2 and 3, respectively. This is due in large part to the separation between intervention and outcome by a horizon of four time steps, which introduces substantial uncertainty and functional complexity into the process. Across tasks, we see that no method achieves the best results consistently: The best ATE estimator is not the best CATE estimator, the best factual fit (e.g., S-learner (RF) in Task 1) does not always lead to the best CATE estimates, and the best CATE estimator in Tasks 1–2 is not the best estimator in Task 3. However, S-learner (XGB) consistently achieves the best (Task 3) or next-best (Tasks 1–2, CATE and ATE) results. This points to the benefit of constructing several evaluation tasks for understanding which methods are preferable. For example, in Figure 2, we see that T-learner (Ridge) consistently overestimates the CATE while other methods underestimate it, albeit less severely. This is likely caused by compounding bias introduced by combining two regularized estimators in the T-learner.
Table 3: Results estimating CATE on IncomeSCM-1.0.CATE, conditioned only on hours-per-week (HPW) with the minimal adjustment set of Task 2. $R^{2}$ and 95% CI computed from 1000 bootstrap samples uses CATE estimated from counterfactual samples within each HPW bin as ground truth.
Estimator $R^{2}$ CATEHPW ($\uparrow$)
S-learner (XGB) 0.40 (-0.32, 0.77)
S-learner (RF) -0.14 (-1.34, 0.54)
T-learner (Ridge) -1.85 (-3.95, -0.70)
T-learner (XGB) 0.06 (-0.84, 0.58)
T-learner (RF) -0.12 (-1.27, 0.57)
Figure 2: Estimated CATE conditioned on hours-per-week, computed by stratifying CATE estimates w.r.t. the minimal conditioning set of Task 2.
@@ 6 Discussion
We have presented the simulator and structural causal model IncomeSCM-1.0, modeled on and fit partially to the Adult data set, to create benchmark tasks for causal effect estimation. The simulator exemplifies a general strategy for creating simulators with causal structure from cross-sectional data sets with challenging complexity. Our benchmarking results show that models with comparable fit to factual outcomes can vary greatly in their estimated causal effects, highlighting the challenging nature of the tasks and the importance of developing purpose-built causal estimators.
The current simulator and benchmark data set have several limitations. First, all trajectories are equally long, with a fixed horizon of $T$ years. This matches observational studies with a fixed follow-up window but does not represent a common challenge: censoring. Specifically, there are no subjects who leave the “study”, either by choice or due to death. Since the age distribution of the initial state follows the census data behind the Adult data set, if the horizon $T$ is set to 100, the simulator would generate samples of humans well over 150 years of age. Similarly, missing values are a challenge in many observational studies but are not generated by the simulator in its current form. Studying the effects of missingness on the ability of estimators to adjust for confounding bias is an interesting direction for future work. Finally, the estimators used to fit initial-state mechanisms were fit with a single set of hyperparameters. It is likely that the fit to Adult could be improved slightly by optimizing these. However, achieving optimal fit is not the purpose of this work.
@@ Appendix A Additional information
    The data is stored as .pkl files of Pandas data frames (created by df.to_pickle(...)). The dataframes can be loaded into Python using Pandas.read_pickle(...)

### v2

@@ 5 Experiments
We see a generally good match in the base statistics of variables, with the biggest discrepancy for income>50k. This is due to renormalizing the simulator samples to achieve an average income of \$70 000 before applying noise and correcting for the effects of studies and workclass = Without-pay. The discriminative performance for the fitted samplers (AUC/$R^{2}$) varies greatly between variables, some (such as marital-status) being more predictable from their parents in the posited causal graph than others (such as capital-net). This is expected since the original data set was not constructed with predicting all variables in mind, nor are all available variables used for prediction, only the causal parents. However, it does mean that the values of variables with low discriminative performance are largely determined by exogenous noise.
Table 2: Results estimating CATE and ATE using the full IncomeSCM-1.0.CATE data set. $R^{2}$ is the coefficient of determination, AE is the absolute error. AUC/$R^{2}$ CV are the (non-random) 5-fold cross-validation AUC of the propensity estimate, and $R^{2}$ of the observed outcome, for propensity and regression estimators, respectively. For $R^{2}$ CATE and AE ATE, the confidence intervals ($\alpha=0.05$) are computed by 1000 bootstrap iterations over test samples. An unbiased estimate of the ATE is \$40 919. Empty cells are intentionally left out as they don’t apply to the corresponding estimators.
@@
“Full” adjustment set: All covariates
IPW (LR) 18645 (18121, 19166) 0.98
IPW (RF) 57025 (56501, 57546) 0.98
IPW-W (LR) 4878 (4354, 5398) 0.98
IPW-W (RF) 17406 (16882, 17927) 0.98
Match (EU-NN) 7147 (6623, 7668) -1.56
S-learner (Ridge) -0.00 (-0.00, -0.00) 400 (22, 901) 0.79
S-learner (XGB) 0.15 (0.14, 0.15) 533 (90, 1023) 0.84
S-learner (RF) 0.03 (0.03, 0.03) 487 (50, 985) 0.84
T-learner (Ridge) 0.24 (0.23, 0.25) 1696 (1213, 2164) 0.80
T-learner (XGB) 0.24 (0.23, 0.25) 207 (10, 574) 0.83
T-learner (RF) 0.30 (0.29, 0.31) 2719 (2256, 3165) 0.84
“Minimal” adjustment set: Direct causes of treatment only
IPW (LR) 12106 (11582, 12627) 0.97
IPW (RF) 58118 (57594, 58638) 0.98
IPW-W (LR) 3369 (2845, 3890) 0.97
IPW-W (RF) 16133 (15609, 16653) 0.98
Match (EU-NN) 8937 (8413, 9458) -1.56
S-learner (Ridge) -0.00 (-0.00, -0.00) 1701 (1177, $2$) 0.78
S-learner (XGB) 0.14 (0.14, 0.14) 1528 (1020, 2012) 0.80
S-learner (RF) 0.02 (0.02, 0.03) 1143 (630, 1643) 0.80
T-learner (Ridge) 0.20 (0.18, 0.21) 6104 (5630, 6595) 0.79
T-learner (XGB) -0.13 (-0.15, -0.11) 16549 (16012, 17095) 0.79
T-learner (RF) 0.16 (0.14, 0.17) 10931 (10453, 11406) 0.80
@@
We include regression estimators in the form of T-learners, which fit one model per potential outcome, and S-learners, which treat the intervention as an input to a single model, mixed in with adjustment variables (Künzel et al., 2019). For both, we compare multiple base estimators in ridge regression (Ridge), random forests (RF), and XGBoost (XGB). Second, we use two classical inverse-propensity weighting estimators, the Horvitz-Thompson (IPW) and Hayek or “Weighted” (IPW-W). Finally, we include a nearest-neighbor matching estimator with a Euclidean metric. The IPW, matching and Ridge-regression S-learner estimators return estimates only of the ATE, not the CATE. For a more detailed description of all estimators, see Appendix E.
Hyperparameters for all estimators (see Appendix Table 6) were selected based on 5-fold cross-validation of 20 uniformly sampled hyperparameter settings. The estimators were then refitted to the entirety of the observational training data. Producing all results, including fitting the simulator, sampling, estimating and evaluating causal effects, took less than 1 hour on a 2023 M2 MacBook Pro. The results were replicated on a 2 x 16 core Intel(R) Xeon(R) Gold 6226R CPU @ 2.90GHz with minimal differences except for XGBoost which is known to yield different fits for different machines.
@@ Results.
We present the results for causal effect estimation Tasks 1 & 2 in Table 2 and for Task 3 in Table 3 and Figure 2. IPW methods and S-learner (Ridge) are excluded from the latter two since these methods do not return heterogeneous effect estimates, only estimates of the ATE. The fit to observables was good for both propensity models (AUC $\approx 0.98$ in predicting observed treatments) and regression estimators ($R^{2}\approx$ 0.80 in predicting observed outcomes). Despite this, all causal estimation tasks were challenging for the methods to solve. The best methods achieved CATE $R^{2}$ of only 0.30, 0.20 and 0.33 on Tasks 1, 2 and 3, respectively, falling short of explaining the majority of variation in these effects with the respective conditioning set. This is likely in large part to the separation between intervention and outcome by a horizon of six time steps (years), which introduces substantial uncertainty and functional complexity into the process.
No method is consistently at the top: The best CATE estimator is not the best ATE estimator and the best CATE estimators are different for Tasks 1, 2 and 3. More importantly, the best factual fit (e.g., S-learner (RF) in Task 1) does not yield the best effect estimates. Despite great discrimination of the fitted propensity scores, IPW estimators generally returned ATE estimates with large error. These results show the importance of constructing several evaluation tasks to understand which methods are preferable in which circumstances. For example, in Figure 2, we see that S-learner (XGB) captures very little of the variation in CATE with education, despite passable performance in Task 1.
Table 3: Results estimating CATE on IncomeSCM-1.0.CATE, conditioned only on education with the full adjustment set of Task 1. $R^{2}$ and 95% CI computed from 1000 bootstrap samples uses CATE estimated from counterfactual samples within each education bin as ground truth.
Estimator $R^{2}$ CATEEDU ($\uparrow$)
S-learner (XGB) 0.33 (0.21, 0.44)
S-learner (RF) -0.02 (-0.12, 0.07)
T-learner (Ridge) 0.02 (-0.30, 0.27)
T-learner (XGB) -0.36 (-0.75, -0.04)
T-learner (RF) -0.61 (-1.25, -0.15)
Figure 2: Estimated CATE conditioned on education (numeric) by stratifying CATE estimates w.r.t. the full conditioning set of Task 1.
@@ 6 Discussion
We have presented the simulator and sequential structural causal model IncomeSCM-1.0, modeled on the Adult data set. The simulator exemplifies a general strategy for creating simulators with causal structure from cross-sectional data sets and use these to create benchmarks for causal effect estimation with challenging complexity. However, the current version and benchmark data set have several limitations. First, all trajectories are equally long, with a fixed horizon of $T$ years. This matches observational studies with a fixed follow-up window but is not representative of a common data analysis challenge: censoring. Specifically, no subjects leave the “study”, either by choice or due to death. Since the age distribution of the initial state follows the census data behind the Adult data set, if the horizon $T$ is set to 100, the simulator would generate samples of humans well over 150 years of age. Similarly, missing values are a challenge in many observational studies but are not generated by the simulator in its current form. Studying the effects of missingness on the ability of estimators to adjust for confounding bias is an interesting direction for future work. Finally, the estimators used to fit initial-state mechanisms were fit with a single set of hyperparameters. It is likely that the fit to Adult could be improved slightly by optimizing these.
Our benchmarking results show that models with comparable fit to observed variables can vary greatly in the quality of their causal effects estimates, highlighting the challenge of the tasks and the importance of developing purpose-built data sets, causal estimators and model selection criteria.
@@ Appendix A Additional information
    The data is stored as .pkl and .csv files of Pandas data frames (created by df.to_pickle(...) and df.to_csv(...)). The dataframes can be loaded into Python using Pandas.read_pickle(...) or Pandas.read_csv(...)
@@
  8. 8.
    Metadata in the Croissant format is included in https://github.com/Healthy-AI/IncomeSCM/blob/main/metadata.json
@@ Appendix E Estimator details
Regression adjustment estimators directly model the outcome $Y$ as a function of the intervention $A$ and adjustment variables $Z$ which guarantee identifiability. Among these, we include T-learners, which fit one model per potential outcome (intervention), and S-learners, which treat the intervention as an input to a single model, mixed in with adjustment variables (Künzel et al., 2019). T-learners use a different base estimator $\hat{mu}_{a}(z)$ to model the potential outcome $Y(a)$ of each intervention $a$, as a function of a chosen adjustment set $Z$. The CATE is then estimated as
$\hat{\mbox{CATE}}(z)=\hat{\mu}_{1}(z)-\hat{\mu}_{0}(z).$
S-learners fit a single base estimator $\hat{\mu}(z,a)$ as a function of both the adjustment set and the intervention. They estimate the CATE as
$\hat{\mbox{CATE}}(z)=\hat{\mu}(z,1)-\hat{\mu}(z,0).$
and both estimate the ATE for a set of $n$ samples with adjustment set covariates $z_{i}$ as
$\hat{\mbox{ATE}}=\frac{1}{n}\sum_{i=1}^{n}\hat{\mbox{CATE}}(z_{i})$
Inverse propensity-weighting (IPW) estimators model the intervention variable instead, as a function of the adjustment set, and re-weigh outcomes to estimate treatment effects. We use two classical estimators, the Horvitz-Thompson (IPW) and Hayek or “Weighted” (IPW-W) estimators of the average treatment effect (ATE). Both fit a function $\hat{e}(z)\approx p(A=1\mid Z=z)$ which is used to re-weight observed outcomes to form an estimate of the potential outcomes and the causal effect. Let $a_{i}\in\{0,1\}$ indicate the binary intervention for subject $i$, $z_{i}$ their covariates from the adjustment set and $y_{i}$ the observed outcome corresponding to $a_{i}$.
$\hat{\mu}_{0}$ $=\frac{1}{n}\sum_{i=1}^{n}w^{0}_{i}y_{i}$ (1)
$\hat{\mu}_{1}$ $=\frac{1}{n}\sum_{i=1}^{n}w^{1}_{i}y_{i}$ (2)
where
$w^{0}_{i}=\frac{1-a_{i}}{1-e(z_{i})}\;\;\;\;w^{1}_{i}=\frac{a_{i}}{e(z_{i})}$
for the Horvitz-Thompson estimator and
$w^{0}_{i}=\frac{1-a_{j}}{1-e(z_{j})}/\left(\sum_{j}\frac{1-a_{j}}{1-e(z_{j})}\right)\;\;\;\;w^{1}_{i}=\frac{a_{i}}{e(z_{i})}/\left(\sum_{j}\frac{a_{j}}{e(z_{j})}\right)~{}.$
for the Hayek estimator. The ATE is then estimated as
$\hat{\mbox{ATE}}=\hat{\mu}_{1}-\hat{\mu}_{0}.$
Finally, we include a nearest-neighbor matching estimator with a Euclidean metric. The estimator computes the nearest neighbor $j(i)$ for all samples $i\in[n]$ according to the chosen metric and estimates,
$\hat{\mu}_{0}$ $=\frac{1}{n}\sum_{i=1}^{n}\left[(1-a_{i})y_{i}+a_{i}y_{j(i)}\right]$ (3)
$\hat{\mu}_{1}$ $=\frac{1}{n}\sum_{i=1}^{n}\left[a_{i}y_{i}+(1-a_{i})y_{j(i)}\right]$ (4)
and
$\hat{\mbox{ATE}}=\hat{\mu}_{1}-\hat{\mu}_{0}.$
For Tasks 1–2, all estimators use the same conditioning set as the adjustment set. For Task 3, the CATE is estimated as in Task 1 but stratified and aggregated in 16 numeric bins of education. Ground truth for each task is constructed analogously but with sampled potential outcomes in place of $\hat{\mu}$.